A candy store has two types of jelly beans, one selling for $2 per pound and the other selling for $4 per pound. The beans are to be mixed to provide 100 lb of a mixture selling for $2.50 per pound. How much of each type of jelly bean should be used to form 100 lb of the mixture?
Answer by mananth(16946)

A candy store has two types of jelly beans, one selling for $2 per pound and the other selling for $4 per pound. The beans are to be mixed to provide 100 lb of a mixture selling for $2.50 per pound. How much of each type of jelly bean should be used to form 100 lb of the mixture?
...
$2 per pound ------------x pounds
$4 per pound. -----------100-x
mixture 100 lb $2.50 per pound.
...
2x+4(100-x)=100*2.50
2x+400-4x=250
-2x+400=250
-400 -400
-2x=250-400
-2x=-150
/-2
x=75 pounds of $2 beans
25 pounds of $4 beans
...
m.ananth@hotmail.ca